OpenStudy (unicornsession):

Which expression is equivalent to (-2c/d^2)^3 ? A. (6c^3/d^6) B. (8c^3/d^6) C. (-6c^3/d^5) D. (-8c^3/d^5)

OpenStudy (anonymous):

\[(\frac{ -2c }{ d ^{2} })(\frac{ -2c }{ d ^{2} })(\frac{ -2c }{ d ^{2} })\]

OpenStudy (unicornsession):

umm thx

OpenStudy (anonymous):

Anytime, Multiple top with top and bottom with bottom

OpenStudy (unicornsession):

but i dont know how to like do tht thts y i asked ths Q. @Hydra_Santana

OpenStudy (anonymous):

Alright. I do not want to just give you the answer. What I did was write the "stuff" inside the parenthesis 3 times because the exponent was 3

OpenStudy (anonymous):

so the top for example is, (-2c)(-2c)(-2c) When you multiple it out you get 3 c's in total and a -8, That is, \[-8c ^{3}\]

OpenStudy (unicornsession):

oh i get it now thx

OpenStudy (anonymous):

Great!